OpenStudy (anonymous):

Please Help!! What are the fourth roots of -64??

OpenStudy (phi):

do you want all 4 roots, including the complex roots?

OpenStudy (anonymous):

Yes, please!

OpenStudy (phi):

one way is write this in polar coordinates the fourth root of -64 is \[ \left(64 e^{i\pi}\right)^{\frac{1}{4}} = (64)^{\frac{1}{4}} e^{\frac{i\pi}{4}} \]

OpenStudy (phi):

the roots are spaced equally around the unit circle

OpenStudy (anonymous):

I am still unsure of this.. the possible answers are in the form: r(cos theta + i sin theta)

OpenStudy (anonymous):

does this make sense?

OpenStudy (anonymous):

What the hell. I am unsure how to find the such a root of a number as it is.

OpenStudy (phi):

remember that \[ r e^{i\theta} = r(\cos(\theta)+i \sin(\theta)) \]

OpenStudy (phi):

64 is 8*8 and sqrt(8) is 2sqrt(2) This will be the magnitude the first root will be at \(\theta= \frac{\pi}{4} \) this is written either as \[ 2 \sqrt{2} e^{i\frac{\pi}{4}} = 2 \sqrt{2} \left( \cos(\frac{\pi}{4})+i \sin(\frac{\pi}{4})\right)\] the 4 roots are spaced equally around \(2\pi\), so their spacing is \[\frac{2\pi}{4}= \frac{\pi}{2}\] so the next root will be at \(\theta= \frac{\pi}{4}+ \frac{\pi}{2} = \frac{3\pi}{4}\) keep adding pi/2 until you get all 4 roots